OpenStudy (akohl):

A jetliner is traveling east from Salt Lake City to Washington DC, a distance of 1,850 miles. The jetliner travels at an average air speed of 590 mph relative to the air. At the jetliner's altitude, the wind is blowing 36 mph to the west. How long does the journey take? @abhisar

OpenStudy (abhisar):

What will be the velocity of jetliner relative to ground ?

OpenStudy (akohl):

626

OpenStudy (abhisar):

Do you know the answer ?

OpenStudy (akohl):

Is it 1850/626

OpenStudy (akohl):

2.96 hr

OpenStudy (abhisar):

This is the given answer ?

OpenStudy (akohl):

That's way I got no given answer

OpenStudy (abhisar):

I think, 590 is the relative velocity wrt air. So 590 = x+36 --> 590-36=554=x=velocity of jetliner wrt to ground.

OpenStudy (abhisar):

Now time=1850/554=3.33 hr

OpenStudy (akohl):

Thanks

OpenStudy (abhisar):

Is it the correct answer ? How will we know ?

OpenStudy (akohl):

We don't know

OpenStudy (abhisar):

When will we know ?

OpenStudy (akohl):

If 590 is relative does that mean with air factored in

OpenStudy (abhisar):

i am not sure about this one. I will be to explain it properly only if i come to know the answer.

OpenStudy (abhisar):

@Abmon98 can u check it ?

OpenStudy (akohl):

K I'll let u know when we go over it in claas

OpenStudy (abmon98):

Average Speed=Total Distance/Total time 590*T=1850 T=1850/590

OpenStudy (abhisar):

Sure ?

OpenStudy (akohl):

Thanks

OpenStudy (abmon98):

i am not sure because of the wind part.

OpenStudy (akohl):

Yea its worded poorly with relative to the wind

OpenStudy (abhisar):

yeah...i m in a same state.. Sincew the given velocity is relative to air. To calculate the the time we need velocity wrt to ground.

OpenStudy (akohl):

I think its 554

OpenStudy (abhisar):

me too !

OpenStudy (abhisar):

but then the next moment i am confused with 626 as u calculated. But more inclined towards 554

OpenStudy (akohl):

I think its 554 bc the air sees it as 590 bc it goes 36 in the opposite dorection

OpenStudy (akohl):

Direction

OpenStudy (abhisar):

yeah that's why @akohl

OpenStudy (akohl):

Gtg thanks for ur help ill let u know what my teacher says tommorrow

OpenStudy (abmon98):

is the speed 626 i got 626 as my answer?

OpenStudy (abhisar):

it's not clear. He will tell us tmrw i think. I am also confused.

OpenStudy (abhisar):

Let's see...btw thanx @Abmon98 !

OpenStudy (abmon98):

your welcome :)

OpenStudy (abmon98):

 Excuse my drawing XD

OpenStudy (abhisar):

whoa !! That's fab..

OpenStudy (abhisar):



OpenStudy (abhisar):

excuse me for my no drawing XD

OpenStudy (abmon98):

lol, you have got a good point wind is travelling against the air planes direction but the problem here is that that the destination is towards the east so when i drew a displacement vector i couldn't find a way to join them together other that.

OpenStudy (abhisar):

:D